### v1

@@ 4.4 Comparison with State-of-the-art Methods
[…]
CSVR(Ours) C3D UCF101 120 $112\times 112$ 16 88.2 52.2
CSVR(Ours) S3D UCF101 120 $112\times 112$ 16 88.0 54.8
CSVR(Ours) R(2+1)D UCF101 120 $112\times 112$ 16 90.3 56.5
@@
Spation-Temp[9] C3D Kinetics400 20 $112\times 112$ 32 61.2 33.4
DCM[160] C3D Kinetics400 120 $112\times 112$ 16 83.4 52.9
CoCLR[155] S3D Kinetics400 500 $128\times 128$ 32 87.9 54.6
VideoMoCo[14] R(2+1)D Kinetics400 200 $112\times 112$ 32 78.7 49.2
DCM[160] R(2+1)D Kinetics400 120 $112\times 112$ 16 85.7 54.0
FAME[153] R(2+1)D Kinetics400 200 $112\times 112$ 16 84.8 53.5
VCL[3] R(2+1)D Kinetics400 100 $112\times 112$ 16 86.1 54.8
TCLR[69] R3D Kinetics400 100 $112\times 112$ 16 84.1 53.6
TE[18] R3D Kinetics400 200 $128\times 128$ 16 87.1 63.6
DynamoNet[162] STCNet* Kinetics400 200 $112\times 112$ 16 88.1 59.9
[…]
MSCL[1] R3D Kinetics400 400 $112\times 112$ 16 90.7 62.3
CSVR(Ours) C3D Kinetics400 120 $112\times 112$ 16 89.8 60.8
CSVR(Ours) S3D Kinetics400 120 $112\times 112$ 16 88.2 59.2
CSVR(Ours) R(2+1)D Kinetics400 120 $112\times 112$ 16 90.8 63.0
CSVR(Ours) R3D Kinetics400 200 $112\times 112$ 16 92.5 64.5
Table 8: Video Retrieval. Comparison of CSVR with the state-of-the-art on nearest-neighbour video retrieval on UCF101 and HMDB51. Given query test clips, our goal is to find training clips that are from the same class using Recall at $k$(R@$k$) metric.
@@
CSVR(Ours) R(2+1)D 120 64.4 80.2 87.8 93.1 34.1 57.2 71.0 81.4
@@
Action Recognition. We firstly compare CSVR with the previous self-supervised video representation learning methods in terms of the downstream action recognition task. As shown in Table 7, when pretraining on UCF101, our CSVR consistently outperforms all existing methods with the same backbone. Moreover, when pretraining on large-scale dataset Kinetics400, our method also achieves the best results of $92.5\%$ and $64.5\%$ on UCF101 and HMDB51, respectively. We can also observe that the trained model by our CSVR surpasses the methods which use other modality of data like optical flow [155], multi-modal information [160]. Besides, our CSVR needs fewer pretraining epochs than most existing methods. All these results demostrate that CSVR is both effective and efficient for self-supervised video representation learning on action recognition.
Video Retrieval. Besides action recognition, we also conduct evaluations for the downstream video retrieval task on two human action datasets UCF101 and HMDB51. Table 8 shows the results of video retrieval of each method in terms of R@K, where R@K denotes the recall rate of top-K. All methods are pretrained on Kinetics400 with the resolution of 112$\times$112 for fair comparisons. The similar conclusion can be obtained that our CSVR achieves the best results on both UCF101 and HMDB51, showcasing its superiority for self-supervised features learning for human action understanding.
5 Conclusion

### v2

@@ IV-D Comparison with State-of-the-art Methods
Vi2CLR[161] S3D UCF101 300 $112\times 112$ 16 82.8 52.9
CSVR(Ours) C3D UCF101 200 $112\times 112$ 16 88.3 52.3
CSVR(Ours) S3D UCF101 200 $112\times 112$ 16 88.1 54.9
CSVR(Ours) R(2+1)D UCF101 200 $112\times 112$ 16 90.4 56.6
@@
DCM[160] C3D Kinetics-400 120 $112\times 112$ 16 83.4 52.9
CoCLR[155] S3D Kinetics-400 500 $128\times 128$ 32 87.9 54.6
VideoMoCo[14] R(2+1)D Kinetics-400 200 $112\times 112$ 32 78.7 49.2
DCM[160] R(2+1)D Kinetics-400 120 $112\times 112$ 16 85.7 54.0
FAME[153] R(2+1)D Kinetics-400 200 $112\times 112$ 16 84.8 53.5
VCL[3] R(2+1)D Kinetics-400 100 $112\times 112$ 16 86.1 54.8
TCLR[69] R3D Kinetics-400 100 $112\times 112$ 16 84.1 53.6
TE[18] R3D Kinetics-400 200 $128\times 128$ 16 87.1 63.6
DynamoNet[162] STCNet* Kinetics-400 200 $112\times 112$ 16 88.1 59.9
Vi2CLR[161] S3D Kinetics-400 300 $112\times 112$ 16 89.1 55.7
MSCL[1] R3D Kinetics-400 400 $112\times 112$ 16 90.7 62.3
KnowledgeGuide[167] C3D Kinetics-400 400 $112\times 112$ 16 82.7 47.0
KnowledgeGuide[167] R(2+1)D Kinetics-400 400 $112\times 112$ 16 82.4 46.7
CSVR(Ours) C3D Kinetics-400 200 $112\times 112$ 16 92.8 62.8
CSVR(Ours) S3D Kinetics-400 200 $112\times 112$ 16 90.2 61.2
CSVR(Ours) R(2+1)D Kinetics-400 200 $112\times 112$ 16 92.8 64.0
CSVR(Ours) R3D Kinetics-400 200 $112\times 112$ 16 94.5 65.5
Action Recognition on UCF101 and HMDB51. We firstly compare CSVR with the previous self-supervised video representation learning methods in terms of the downstream action recognition task. As shown in Table IV-D, when pretraining on UCF101, our CSVR consistently outperforms all existing methods with the same backbone. Moreover, when pretraining on large-scale dataset Kinetics-400, our method also achieves the best results of $94.5\%$ and $65.5\%$ on UCF101 and HMDB51, respectively. We can also observe that the trained model by our CSVR surpasses the methods which use other modality of data like optical flow [155], multi-modal information [160]. Besides, our CSVR needs fewer pretraining epochs than most existing methods. All these results demostrate that CSVR is both effective and efficient for self-supervised video representation learning on action recognition.
TABLE VII: Action Recognition on SSv2. Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task.
{tabu}
l—ccccc—cc Method Backbone Pretrain Dataset Epochs Resolution Clip Length Top-1
VideoMoCo[14] R(2+1)D Kinetics-400 200 $112\times 112$ 32 19.5
$\rho$MoCo[65] SlowFast-R50 Kinetics-400 200 $224\times 224$ 16 54.4
CORP[166] I3D Kinetics-400 800 N/A 16 48.8
SVT[165] ViT-B Kinetics-400 20 $224\times 224$ 16 59.2
CSVR(Ours) R(2+1)D Kinetics-400 200 $112\times 112$ 16 59.2
Action Recognition on SSv2. Comparing to UCF101 and HMDB51, SSv2 exhibits less scene bias than previous datasets and require a stronger temporal understanding to accurately classify actions. As shown in Table. IV-D, our method outperforms the best previous methods using 3D-CNN backbones by an absolute margin of 4.8% on SSv2 when pretrained on Kinetics-400. In addition, the performance of CSVR on the SSv2 action recognition task is comparable to that of SVT[165], which uses a stronger ViT backbone[89] and higher video resolution. Such results demonstrating its suitability for video datasets captured in controlled, real-world settings.
TABLE VIII: Video Retrieval. Comparison of CSVR with the state-of-the-art on nearest-neighbour video retrieval on UCF101 and HMDB51. Given query test clips, our goal is to find training clips that are from the same class using Recall at $k$(R@$k$) metric.
@@
KnowledgeGuide[167] R(2+1)D 400 52.6 70.1 77.4 85.2 22.2 45.9 59.9 74.2
KnowledgeGuide[167] C3D 400 49.9 68.3 77.4 84.9 22.4 46.9 60.3 73.1
CSVR(Ours) R(2+1)D 200 64.4 80.2 87.8 93.1 34.1 57.2 71.0 81.4
@@
Video Retrieval. Besides action recognition, we also conduct evaluations for the downstream video retrieval task on two human action datasets UCF101 and HMDB51. Table VIII shows the results of video retrieval of each method in terms of R@K, where R@K denotes the recall rate of top-K. All methods are pretrained on Kinetics-400 with the resolution of 112$\times$112 for fair comparisons. The similar conclusion can be obtained that our CSVR achieves the best results on both UCF101 and HMDB51, showcasing its superiority for self-supervised features learning for human action understanding.
TABLE IX: Computing cost comparing with popular self-supervised video representation learning methods. All methods are pretrained on K400.
{tabu}
l—ccc—c Method Backbone GFLOPsParameters UCF Top-1
MoCo[124] C3D 38.7 33.4M 60.5
CoCLR[155] S3D 36.0 9.1M 87.9
SpeedNet[98] S3D 36.5 9.1M 81.1
$\rho$MoCo[65] SlowFast-R50 83.5 31.8M 92.8
CSVR C3D 40.4 33.4M 92.8
CSVR S3D 37.7 9.1M 90.2
CSVR R(2+1)D 40.4 15.4M 92.8
Computational cost. To compare computational costs, we enumerate the model parameters of some popular self-supervised video representation learning methods pretrained on Kinetics-400 in Table. IV-D. Following the common practices in [65, 168, 87], we report the parameters of each model during the inference phase. As shown, CSVR significantly improves accuracy in terms of the UCF101 action recognition downstream task compared to methods using the same backbone. Compared to MoCo[124], which also uses C3D as the backbone, our method boosts performance by over 30% on the UCF101 dataset. Additionally, CSVR using R(2+1)D as the backbone achieves comparable performance with $\rho$MoCo[65], which uses the heavier SlowFast [33] as the backbone. Noting that, CSVR has higher GFLOPs compared to the methods using the same backbone since we have an additional generative pose prediction branch.
V Limitation
CSVR is a self-supervised video representation learning framework with multiple branches. These branches independently learn dynamic and static features, which are then optimized collaboratively. To manage the complexity of optimizing multiple neural networks during pretraining, we opt for lightweight CNNs instead of stronger, but larger backbones like Transformers. Additionally, we see potential for integrating more advanced generative methods (e.g., diffusion models[171]) and contrastive learning approaches (e.g., BYOL[65]) into our framework, given that the primary goal of the dynamic and static branches is to learn comprehensive video representations. Exploring these enhancements further will be a focus of our future work.
VI Conclusion